Minutes of nanoBPM meeting on 19 September, 2003

"Nano BPM collaboration plan", T. Tauchi

(transparencies, 6 pages, pdf, 637KB )

After Marc's talk in the video meeting, we organized the nano-BPM collaboration in following three items@in a sequential order.

  1. Establishment of nano-meter resolution
  2. Stabilization of two isolated systems ( Phase I of Marc's talk)
  3. Fast feedforward/feedback tests ( Phase II of Marc's talk)
First task is a construction of 3-cavityBPM systems at KEK and also at SLAC/LLNL. SLAC/LLNL system was already under construction with existing cavityBPMs, while KEK system is still at designing stage. At KEK, prototypes of Y' and integrated X and X' movers will be manufactured in order to verify a specific hinge mechanism controlled by Piezo devices. The KEK system consists of 3 mover systems, 3 cavityBPMs and a reference system, which is mounted on a girder (granite table). The movers must accept three control singnals of the reference system, inertia sensors and optical anchors, at least. If budgets are available, major part of the KEK system will be constructed in this fiscal year. Therewithal, um feed-forward will be tested in order to stabilize the ATF beam vertically within +/- a few um at the extraction line.

For the stabilization between two systems, there must be three major schemes;

The "remote" may conceptually contradict the "passive" unless the optical anchor is also set on the girder. Anyway, the um feed-forwarded ATF beam is essential for a reference line.

The final experimental phase is a demonstration of beam control at nano-meter level with combination with FEATHER or FONT system, that is the fast feedforward/feedback test in a bunch-train. In this phase, the multi-train extraction must be very important.

He finally showed a possible layout; the SLAC/LLNL system at QD7-QF6, the KEK system at QD8-QF7 and the FEATHER/FONT at QF6-QD8. The optical anchor may be set at polarized positron area.

"GM measurement at ATF-floor and the reference system", H.Yamaoka

(transparencies, 21 pages, pdf, 2.6MB , or power point file, 9.6MB, and 3 figures of integrated amplitudes are here(PDF,619KB).)

Collaborating with Sugahara-san and Masuzawa-san, we (Yamaoka, Nicolas and Tauchi) measured ground motions at ATF. Three sensors (NS, EW, UD) were set each at two (floor) points in vicinity of QD7-QF6 and QF6-QD8, at north side of the extraction line. The sensor is an accelerometer (MG-102 of Tokyo-kiki) with resolution of 10-6 G ( about 1 mgal) . Sampling frequency was 256Hz and sensitive frequency region is greater than 1 Hz. Conditions are listed below;

Every data contain 78sec periodical distinct "electrical" noises whose origin was unknown, although the noises should have no effect on present results. There was a clear difference in power spectra between day (1pm) and night (10pm). A prominent peak appeared at 3 Hz only in day. The day spectrum is very similar to that measured at Oho experimental hall at 10am. Good coherency(~1) was observed around 3 Hz and fair one(>0.5) was observed below 10Hz.

Daily variation from 30nm(night) to 80nm(day) was observed in the integrated spectra of f >1Hz (8/26-27,8/29-30) while the 5 and 10Hz integrated ones have negligibly small variation. In Sunday, the daily variation became very small and the 1Hz integrated amplitudes were around 30nm. On Monday morning at 7am, 9/1, large ground motions were observed around 1 Hz, which were about 3 times larger than usual ones. They have very good coherency (~1) in all directions. On 8/28, two of three water lines were stopped for maintenance from 13:30 to 15:30. However, no apparent effect was observed in power spectra except for the last 10 minutes (15:20-15:30).

He demonstrated ground motions of the ATF floor by animation. He also updated the spectrum analysis of the reference frame(system) with a ground motion model based on the spectrum at 0:00pm, 1st September. Without using isolators, the 1st resonant peaks appear at 80~90Hz and ~270Hz in X,Z and Y directions, respectively. The peak amplitudes were less than a nano-meter, while typical amplitude(RMS) was about 5 nm at f=1Hz.

"Prototype designs", Y.Honda

Honda explained the design of (X,X') prototype mover. The X' has an inner plate of 6cm thick, which is supported with 8 hinges by outer frame. These hinges can support the plate with 500N (~50kg-G) force. Since the "rotational" center of the plate is not fixed, he devised a hinge structure (e.g. 2.6mm thick and 10mm long with necks at both ends) for enough rigidness so that the center movement can be limited within 1mm. Necessary rotation of 400urad can be achieved by 40um distortion of hinges. Major R&D item would be stability of the rotational center for such distortion. The X and Y movers use 4 simple hinges of 0.3mm thickness and 10mm length, which can accept a distortion of 250um.

He tested Piezo devices with stacked structure, whose dynamic ranges are 80um/140V and 20um/27V in catalog. Resultant values were 17um/30V and 85um/50V with no-load.

He will complete designs of the prototype within this week (until 9/20). The prototypes will be manufactured at the KEK mechanical engineering center.

Next meeting will be held in second week of October.